40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5268     Entry time: Fri Aug 19 09:14:56 2011
Author: steve 
Type: Update 
Category: IOO 
Subject: degrading laser power at atm 

Light into the MC is 20 mW at atm, MC_Transmitted ~10 MW = 400 count

The PMC_T is OK but something else is drifting.

Attachment 1: power@atm.jpg  90 kB  | Hide | Hide all
power@atm.jpg
ELOG V3.1.3-